jackie:

The square of a number is 12 less than 7 times the number. What is the number? –4 or –3 –4 or 3 –3 or 4 3 or 4

Vocaloid:

let "a number" be n "12 less" implies n - 12 "7 times" implies 7n "square of a number" implies n^2 7n - 12 = n^2 solve for n

EvieSwan2405:

The answer must be x=4, because 4*(4-1) = 4*3 =12. And, if negative numbers are allowed, x = -3, because (-3)(-4) = 12. That;s all such numbers. I looked it up thats what I found